# A Study of Plasticity Loss in On-Policy Deep Reinforcement Learning

### v1

@@ 4.3 Architectural interventions do not fully address plasticity loss
Figure 4: Performance of intervention methods compared to baselines of warm-start and reset-all on the gridworld environment. Mean reward is normalized by performance at end of first round. Top row: Training performance. Bottom row: Test performance.
@@
Recently a number of strategies have been introduced to specifically address plasticity loss via changes to the network architecture. These methods were previously validated in the off-policy setting only. Here we consider the behavior of two of these methods as it applies to the on-policy setting: CReLU (Abbas et al.,, 2023) and Plasticity injection (Nikishin et al.,, 2023). We find that both methods fail to address plasticity loss in the permute shift condition. The CReLU method is able to address plasticity loss in the window and expand settings only, while the plasticity inject method underperforms even the warm-start baseline in all three contexts (Figure 4).
4.4 Regularization methods do address plasticity loss
@@
We next consider the class of methods which regularize the weights of the network, either continuously or intermittently. These methods include final layer reset (Nikishin et al.,, 2022), shrink+perturb (Ash and Adams,, 2020), a continuous “soft” variant of shrink+perturb (Dohare et al., 2023b, ), L2 regularization, ReDo (Sokar et al.,, 2023), and regenerative regularization (Kumar et al.,, 2023). We find that the final-layer reset fails to address plasticity loss in any of the three conditions. This is likely because the learned penultimate layer representation is insufficient for mapping states to optimal actions after a shift in state distribution. In contrast, the other five methods mitigate both plasticity loss and the warm-start problem to some extent in all three shift conditions (Figure 4). Of these methods, ReDo performs the worst as it is unable to benefit from the positive transfer which is possible in the cases of the “window” and “expand” conditions, as methods that retain information from previous rounds effectively have access to more training data in these conditions. ReDo underperforms even the warm-start baseline in the “expand” condition.
@@ 4.5 LayerNorm addresses training plasticity loss but not generalization
We also consider the effect of LayerNorm on plasticity loss. It was previously demonstrated in the off-policy setting that LayerNorm is able to mitigate some effects of plasticity loss (Lyle et al.,, 2023). Given that LayerNorm tends to improve learning across a number of contexts (Ba et al.,, 2016), it is important to attempt to separate the effects of LayerNorm on performance generally from its effect on plasticity specifically. We find that LayerNorm resolves plasticity loss in terms of training performance, but is inconsistent in its effect on generalization performance. For example, the LayerNorm model significantly underperforms even the warm-start baseline on the test data of the permute shift condition. We find however that combining soft shrink+perturb or regenerative regularization with LayerNorm results in overcoming both training plasticity loss and deleterious generalization trends (Figure 4).
@@ 4.6 Plasticity loss in CoinRun
We also use the CoinRun environment from the ProcGen suite of tasks to evaluate the set of candidate interventions (Cobbe et al.,, 2019, 2020). ProcGen is built using procedural generation (Cobbe et al.,, 2020), making it possible to study the same three distribution-shift conditions which were considered in the gridworld experiments. Here we present performance results for each of the intervention methods on the CoinRun environment.
Figure 5: Performance of intervention methods compared to baselines of warm-start and reset-all on the CoinRun environment. Mean reward is normalized by performance at end of first round. Top row: Training performance. Bottom row: Test performance.
Overall, we find results largely consistent with earlier results—there is significant plasticity loss present in all three tasks as measured by the gap between the warm-start and reset-all conditions. We also find that reset-final, CReLU, and plasticity injection are all unable to fully resolve the pathology. In contrast, the three most effective methods in the gridworld task (LayerNorm, regenerative regularization, and soft shrink+perturb) are all effective in CoinRun as well. See Figure 5 for plots of the normalized mean rewards over the course of training for all methods. For results on two additional ProcGen environments, see Appendix D.
For round-level performance plots of all methods, see Section D in the appendix. We find that relative efficacy of the intervention methods is consistent across ProcGen environments.
@@ 4.7 Plasticity loss in Montezuma’s Revenge
Montezuma’s Revenge is an Atari game that is often used to benchmark the quality of exploration procedures in RL (Bellemare et al.,, 2013; Salimans and Chen,, 2018). The environment consists of many separate rooms, some only accessible via a locked door. The agent must avoid obstacles, find and use keys, and traverse several rooms before obtaining non-trivial reward. Because of this reward sparsity, exploration is needed to incentivize the agent to advance beyond the periphery of its experience.
Montezuma’s Revenge is designed such that the agent only sees the room it currently occupies. Accordingly, the distribution over input states expands as the agent learns to explore the environment. Unlike other ALE games, this introduction of significantly different states as training progresses induces a level of distribution shift. Within our framework, the distribution shift of Montezuma’s Revenge is most similar to the “expand” condition studied earlier.
@@ 5 Discussion
In contrast, the class of regularization methods, which include L2 regularization, ReDo, shrink+perturb, and regenerative regularization all are able to significantly mitigate plasticity loss in our experiments. Of these methods, soft shrink+perturb displayed the best generalization performance, and can be easily combined with layer normalization. LayerNorm, a now ubiquitous method in supervised learning, was previously introduced to address plasticity loss in the off-policy setting (Lyle et al.,, 2023), and we find that it is also effective in the on-policy setting. Due to its general regularizing effects, LayerNorm increases performance in the absence of distributional shift (Ba et al.,, 2016), suggesting that it is a generally useful method for on-policy reinforcement learning methods such as PPO.
A consistent feature of the class of regularization methods discussed here is that they all normalize the network parameters towards their initial distribution. This has the effect of decreasing the weight magnitude and reducing the number of dead or saturated activation units, as seen in Figure 3. It has recently been proposed that both of these measurements may be a proxy for the underlying curvature of the optimization landscape (Lewandowski et al.,, 2023); studying this connection could be a fruitful avenue for future research.
@@ 6 Related Work
In the context of generalization in continual learning settings, a related issue to plasticity loss is the warm-start problem (Ash and Adams,, 2020). Recent work has suggested that the benefits of methods like shrink+perturb to generalization may depend on the presence of noise in the labels used for training (Zaidi et al.,, 2023). In the absence of label noise, other regularization techniques were found to be competitive with methods that make use of re-initialization.
@@
In the off-policy RL setting the problem of plasticity loss has also been studied under the name of capacity loss (Lyle et al.,, 2022). While plasticity loss refers to a property of the network-task interaction, capacity loss is a more generic term to refer to the properties of a network that are task invariant. This work advocates for a solution similar to that of Kumar et al., (2023), and can thus be considered in the class of methods which regularize the network towards an initialization distribution.
Within the domain of on-policy RL there is a related phenomena to plasticity loss called policy collapse, which refers to a degradation of performance as training progresses, even on a fixed dataset (Dohare et al., 2023c, ). Although these are distinct phenomena, there is likely overlap in the success of mitigation methods. In terms of direct studies of plasticity loss in on-policy RL, both Dohare et al., 2023a and Igl et al., (2020) demonstrate plasticity loss in the on-policy setting, and propose a method to address it.
As part of their analysis, Dohare et al., 2023b propose a plasticity-enhancing method called Continual Backprop, which is another form of regularization towards the initialization distribution, however it is one which is selective on a per-neuron basis. The approach is similar to ReDo (Sokar et al.,, 2023), which we analyze here. It is also related to DrM (Xu et al.,, 2023), which builds on the DrQ algorithm.
It is finally worth mentioning that notions of plasticity loss have been studied in the neuroscience community as well. These range from classic work demonstrating the inability of animal models to fully accommodate new sensory input after developing cognitively without it (Wiesel and Hubel,, 1963), to more recent work characterizing plasticity loss in humans as a causal factor in various psychopathologies (Peled,, 2005; Carhart-Harris et al.,, 2023; Juliani et al.,, 2024).
@@ 7 Conclusion
This work studies the problem of plasticity loss in the context of on-policy deep reinforcement learning. Like in continual learning and off-policy RL, we find plasticity loss is a pervasive issue across a number of different environments and forms of distribution shift. We further find that some methods previously proposed to resolve the issue in other problem settings, such as CReLU and plasticity injection, do not transfer to the on-policy regime. The methods which best resolve plasticity loss are those which act as continual regularizers as opposed to intermittent interventions. Within this class of methods, soft shrink+perturb and regenerative regularization seem to perform best, particularly when used in conjunction with layer normalization.
@@ Appendix B Environment and Model Details
Additional hyperparameters used for Gridworld and CoinRun models are provided in Table 2. Unless indicated otherwise, all experiments involve running five replicates with unique random seeds each. All experiments are conducted using either a single P100 or V100 GPU on a cloud machine.
@@
CoinRun. The CoinRun environment consisted of a 64x64x3 observation space. This was provided to the PPO model as a tensor, and processed by a convolutional encoder with four layers and ReLU activations. Each convolutional layer utilized a kernel size of four and a stride of two. In the permute environment variation, images were broken into 8x8x3 patches and the position of patches was randomized according to a fixed map within each round. A “dual-head” architecture was also used. The action space consisted of the standard nine possible actions provided by the ProcGen environment Cobbe et al., (2020). Episode length varied based on agent behavior. Each round of training consists of 5,000 iterations of inference and model updates. Given ten rounds, the total number of iterations is 50,000.
@@ Appendix D Extended Results
Figure 10: Performance of intervention methods compared to baseline (warm-start) on Gridworld task. Each error bar corresponds to normalized mean episodic reward. Tow row: Training distribution performance. Bottom row: Test distribution performance.
Figure 11: Performance of intervention methods compared to baseline (warm-start) on CoinRun task. Each error bar corresponds to normalized mean episodic reward. Top row: Training distribution performance. Bottom row: Test distribution performance.

### v2

@@ 4.3 Novel architectural methods do not fully address plasticity loss
Figure 4: Performance of intervention methods compared to warm-start and reset-all baselines on the Gridworld environment. Final round mean reward is normalized by the performance at end of the first round, and interval bars denote standard error. Top: Train performance. Bottom: Test performance.
@@
Recently a number of methods have been introduced to specifically address plasticity loss using changes to the network architecture. These methods were previously validated in the off-policy setting only. Here we consider the behavior of two of these methods as it applies to the on-policy setting: CReLU Abbas et al., (2023) and Plasticity injection Nikishin et al., (2023). We find that both methods fail to address plasticity loss in the permute shift condition. The CReLU method is able to address plasticity loss in the window and expand settings only, while the plasticity inject method underperforms even the warm-start baseline in all three contexts (Figure 4).
4.4 Regularization methods address plasticity loss
@@
We next consider the class of methods which regularize the weights of the network, either continuously or intermittently. These methods include final layer reset (Nikishin et al.,, 2022), shrink+perturb (Ash and Adams,, 2020), a continuous “soft” variant of shrink+perturb (Dohare et al., 2023b, ), L2 regularization, ReDo (Sokar et al.,, 2023), and regenerative regularization (Kumar et al.,, 2023). We find final layer reset fails to address plasticity loss in any of the three conditions. In contrast, the other five methods mitigate both plasticity loss and the warm-start problem to some extent in all three shift conditions (Figure 4). Of these methods, ReDo performs the worst as it is unable to benefit from the positive transfer which is possible in the cases of the “window” and “expand” conditions, as methods that retain information from previous rounds effectively have access to more training data in these conditions. ReDo underperforms even the warm-start baseline in the “expand” condition.
@@ 4.5 LayerNorm addresses training plasticity loss but not generalization
Figure 5: Performance of intervention methods compared to warm-start and reset-all baselines on the CoinRun environment. Final round mean reward is normalized by the performance at end of the first round, and interval bars denote standard error. Top: Train performance. Bottom: Test performance.
We also consider the effect of LayerNorm on plasticity loss. It was previously demonstrated in the off-policy setting that LayerNorm was able to mitigate some effects of plasticity loss (Lyle et al.,, 2023). Given that LayerNorm tends to improve learning across a number of contexts (Ba et al.,, 2016), it is important to attempt to separate the effects of LayerNorm on performance generally from its effect on plasticity specifically. We find that LayerNorm resolves plasticity loss in terms of training performance, but is inconsistent in its effect on generalization performance. For example, the LayerNorm model significantly underperforms even the warm-start baseline on the test data of the permute shift condition. We find however that combining soft shrink+perturb or regenerative regularization with LayerNorm results in overcoming both training plasticity loss and deleterious generalization trends (Figure 4).
@@ 4.6 Plasticity loss in CoinRun
We also use the CoinRun environment from the ProcGen suite of tasks to evaluate the set of candidate interventions Cobbe et al., (2019, 2020). ProcGen is built using procedural generation Cobbe et al., (2020), making it possible to study the same three distribution-shift conditions which were considered in the gridworld experiments. Here we present performance results for each of the intervention methods on the CoinRun environment.
Overall, we find results largely consistent with those of the gridworld task. There is significant plasticity loss present in all three tasks as measured by the gap between the warm-start and reset-all conditions. We also find that reset-final, CReLU, and plasticity injection all unable to fully resolve the loss of plasticity. In contrast, the three most effective methods in the gridworld task (LayerNorm, regenerative regularization, and soft shrink+perturb) are all also most effective in CoinRun as well.
See Figure 5 for plots of the normalized mean rewards over the course of training for all methods. For results on two additional ProcGen environments, see Appendix D. For round-level performance plots of all methods, see Section D in the appendix. We find that relative efficacy of the intervention methods is consistent across ProcGen environments.
@@ 4.7 Plasticity loss in Montezuma’s Revenge
Montezuma’s Revenge is an Atari game that is often used to benchmark the quality of exploration procedures in RL Bellemare et al., (2013); Salimans and Chen, (2018). The environment consists of many separate rooms, some only accessible through a locked door. The agent must avoid obstacles, find and use keys, and traverse several rooms before obtaining non-trivial reward. Because of this reward sparsity, exploration is needed to incentivize the agent to advance beyond the periphery of its experience.
Montezuma’s Revenge is designed such that the agent only sees the room it currently occupies. Accordingly, the distribution over input states expands as the agent learns to explore the environment. This introduction of significantly different states induces a level of distribution shift which is unique among ALE environments. Within our framework, the distribution shift of Montezuma’s Revenge is most similar to the “expand” condition studied here.
@@ 5 Discussion
In contrast, the class of regularization methods, which include L2 regularization, ReDo, shrink+perturb, and regenerative regularization all are able to significantly mitigate plasticity loss in our experiments. Of these methods, soft shrink+perturb displayed the best generalization performance, and can be easily combined with layer normalization. LayerNorm, a now ubiquitous method in supervised learning, was previously introduced to address plasticity loss in the off-policy setting Lyle et al., (2023), and we find that it is also effective in the on-policy setting. Due to its general regularizing effects, LayerNorm also increases baseline performance in the absence of distributional shift Ba et al., (2016), suggesting that it is a generally useful method for on-policy reinforcement learning methods such as PPO.
A consistent feature of the class of regularization methods discussed here is that they all normalize the network parameters towards their initial distribution. This has the effect of decreasing the weight magnitude and reducing the number of dead or saturated activation units, as seen in Figure 3. It has recently been proposed that both these metrics may be a proxy for the underlying curvature of the optimization landscape (Lewandowski et al.,, 2023). Studying this connection more deeply would potentially be a fruitful avenue for future research.
@@ 6 Related Work
In the context of generalization in continual learning settings, a related issue to plasticity loss is the warm-start problem (Ash and Adams,, 2020). Recent work has shown that the benefits of methods such as shrink+perturb to generalization were dependent on the presence of noise in the labels used for training (Zaidi et al.,, 2023). In the absence of noisy labels, other regularization methods were found to be competitive with methods that utilize re-initialization such as shrink+perturb.
@@
In the off-policy RL setting the problem of plasticity loss has also been studied under the name of “capacity loss” (Lyle et al.,, 2022). While plasticity loss refers to a property of the network-task interaction, capacity loss is a more generic term to refer to the properties of a network that are task invariant. This work advocates for a solution similar to that of Kumar et al., (2023), and can thus be considered in the class of methods which regularize the network towards an initialization distribution.
Within the domain of on-policy RL there is a related phenomena to plasticity loss called “policy collapse,” which refers to a degradation of performance as training progresses, even on a fixed dataset (Dohare et al., 2023c, ). Although these are distinct phenomena, there is likely overlap in the success of mitigation methods. A novel optimizer strategy “Non-stationary Adam” was proposed to address policy collapse, and although we did not study it, it may provide benefits on the tasks examined here. In terms of direct studies of plasticity loss in on-policy RL, both Dohare et al., 2023a and Igl et al., (2020) show plasticity loss in the on-policy setting, and propose a method to address it.
As part of their analysis, Dohare et al., 2023b propose a plasticity-enhancing method called “Continual Backprop,” which is another form of regularization towards the initialization distribution, however it is one which is selective on a per-neuron basis. This method is similar to ReDo (Sokar et al.,, 2023), which we analyze here. It is also related to DrM (Xu et al.,, 2023), which builds on the DrQ algorithm.
It is finally worth mentioning that notions of plasticity loss have been studied in the neuroscience community as well. These range from classic work demonstrating the inability of animal models to fully accommodate new sensory input after developing cognitively without it (Wiesel and Hubel,, 1963), to more recent work characterizing plasticity loss in humans as a component of various psychopathologies (Peled,, 2005; Carhart-Harris et al.,, 2023; Juliani et al.,, 2024).
@@ 7 Conclusion
In this work we studied the problem of plasticity loss in the context of on-policy deep RL. We find that similar to continual learning and off-policy RL, plasticity loss is an issue across a number of different environments and forms of distributional shift. We find that some methods previously proposed to resolve the issue in other problem settings, such as CReLU and plasticity injection do not transfer to on-policy RL. The methods which best resolve plasticity loss are those which act as continual regularizers as opposed to intermittent interventions. Within this class of methods we find that a soft variant of shrink+perturb combined with LayerNorm performs the best across our evaluated settings, providing a simple and general method for addressing plasticity loss.
@@ Appendix B Environment and Model Details
Additional hyperparameters used for Gridworld and CoinRun models are provided in Table 2. All experiments involve running five repetitions with unique random seeds each. All experiments are conducted using either a single P100 or V100 GPU on a cloud machine. We used preliminary exploratory training runs to determine the necessary number of epochs to ensure convergence within the initial single round of training. In these longer training runs we did not find evidence of policy collapse Dohare et al., 2023c , and thus believe that performance degradation in subsequent rounds can be attributed to plasticity loss. Code which can be used to reproduce our results is available at https://github.com/awjuliani/deep-rl-plasticity.
@@
CoinRun. The CoinRun environment consisted of a 64x64x3 observation space. This was provided to the PPO model as a tensor, and processed by a convolutional encoder with four layers and ReLU activations. Each convolutional layer utilized a kernel size of four and a stride of two. In the permute environment variation, images were broken into 8x8x3 patches and the position of patches was randomized according to a fixed map within each round. A “dual-head” architecture was also utilized. The action space consisted of the standard nine possible actions provided by the ProcGen environment Cobbe et al., (2020). Episode length varied based on agent behavior. Each round of training consists of 5,000 iterations of inference and model updates. Given ten rounds, the total number of iterations is 50,000.
@@ Appendix D Extended Results
Figure 7: Performance of intervention methods compared to baseline (warm-start) on Gridworld task. Shaded region corresponds to normalized mean episodic reward. Tow row: Training distribution performance. Bottom row: Test distribution performance.
Figure 8: Performance of intervention methods compared to baseline (warm-start) on CoinRun task. Shaded region corresponds to normalized mean episodic reward. Top row: Training distribution performance. Bottom row: Test distribution performance.